OpenStudy (anonymous):

tony and edwin each built a rectangular garden. tony's garden is twice as long and twice as wide as edwin's garden. If the area of edwin's garden is 600 square feet, what is the area of tony's garden?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

2400

OpenStudy (anonymous):

tony's garden long 2a , and wide 2b 2a*2b=4ab edwin's garden long a , and wide b a*b=ab Tony's garden bigger 4 times than edwin's so 600*4=2400

OpenStudy (anonymous):

Start with the known facts and translate into equations. First look at the infor on Tony's garden, it is equal to 2(w) and 2(L) as Edwins Garden. Second. Look at Edwins Garden. The area is is 600 sq ft. and we know that Area is the formula A = L(w) so the equation for Edwin is 600 = L(W) If Tonys garden is 2(L) and 2(w) then it is 4 times the size of Edwins. So 600 times 4 equals 2400.

OpenStudy (anonymous):

Thanks!